# On the Potential of Microtubules for Scalable Quantum Computation

### v1

@@ Abstract
We examine the quantum coherence properties of tubulin heterodimers in the Microtubule (MT) lattice. In the cavity-MT model proposed by the authors, according to which the MT interiors are modeled as high-Q quantum-electrodynamics cavities, decoherence-resistant entangled states have been argued to emerge under physiological conditions, with decoherence times of order $\mathcal{O}(10^{-6})$ s. The latter is the result of strong electric-dipole interactions of tubulin dimers with ordered-water dipole quanta in the MT interior. We re-interpret the classical nonlinear (pseudospin) $\sigma$-models, describing the emergent dynamics of solitonic excitations in such systems, as representing quantum coherent (or possibly pointer) states, arising from the incomplete collapse of quantum-coherent dipole states. These solitons mediate dissipation-free energy transfer across the MT networks. We underpin logic-gate-like behavior through MT-associated proteins and detail how these structures may support scalable, ambient-temperature quantum computation, with the fundamental unit of information storage being a quDit associated with the basic unit of the MT honeycomb lattice. We describe in detail the decision-making process, after the action of an external stimulus, during which optimal path selection for energy-loss-free signal and information transport across the MT network emerges. Finally, we propose experimental pathways, including Rabi-splitting spectroscopy and entangled surface plasmon probes, to experimentally validate our predictions for MT-based, scalable quantum computation.
preprint: KCL-PH-TH/2025-18
@@ I Introduction: Solitons, dissipationless energy transfer and quantum entanglement in biosystems
One open issue, relevant to using biomatter for quantum computation, as discussed in [1] in the context of a model for biological microtubules (MT) [2] represented as high-quality quantum-electrodynamics (QED) cavities [3, 4], is the efficiency by which energy (and therefore information) is transduced and transported in biosystems. Specifically, in our works we have argued how the QED cavity model of MT can be essential, and in fact unavoidable, for brain functioning [5], in the sense of transforming the trapped information from sources (stimuli) outside the brain into appropriate electromagnetic signals, which are sent and received by the neurons, resulting in exclusively electromagnetic communication through, e.g. synapses. QED Cavities are known for providing a very efficient way of manipulating quantum entanglement of atoms and photons [6]. In a similar way, we have proposed in [3] that electric dipole quanta of the tubulin dimer walls of the cavity interiors of the MT, are entangled with the dipole quanta in the ordered-water interior, leading to solitonic structures of such dipoles, which can transfer energy and signal in a dissipation-free way. By representing these MT dipole systems in terms of condensed-matter models, e.g. pseudospin non-linear $\sigma$-models as in [7], one can quantify such processes of loss-free energy and signal transduction by means of concrete solitonic solutions of the equations of motion of these models. In our picture, as we discuss in [3, 4, 1] and in this article, such classical solitons may be viewed as a kind of coherent states of the dipole quanta.111 Given that MT are ubiquitous structures in all eukaryotic cells [2], we expect such properties to characterize all living biological matter. For instance, recently it has been made possible to transplant RNA molecules, which are known to carry memory and encode information [8], between snails, by literally injecting one snail’s memories into another snail [9], that is passing down RNA through flesh instead of an intact nervous system. Since living cells prodiuce or modify RNA through the action of external stimuli, our model for information processing may therefore apply to such biological structures as well.
@@
In the 1990’s, a suggestion on the role of quantum effects on brain functioning, and in particular on conscious perception, has been put forward by S. Hameroff and R. Penrose (HP) [15], who concentrated on the microtubules (MT) [2] of the brain cells. Specifically, by considering the tubulin heterodimers conformation as quantum states of a two-state system, they assumed their coherent superposition, which may result in helicoidal solitonic states propagating along the MT,333 Note that the helix is a geodesic path (optimal “minimum distance” path) in the Euclidean cylindrical geometry $S^{1}\otimes\mathbb{R}$, embeddable in $\mathbb{R}^{3}$. and thus being responsible for conscious perception. HP assumed sufficiently long decoherence time for this purpose of order ${\mathcal{O}}(1~{}s)$ in in vivo situations, so that the in vivo system of MT in the brain undergoes, as a result of sufficient growth that allowed it to reach a critical mass/energy, self-collapse related to a quantum gravity environment (orchestrated reduction method), as opposed to the standard environmental non-gravitational decoherence that physical quantum systems are subjected to [13].
@@
$x-x_{0}$ =±$=\pm\int_{\phi(x_{0})}^{\phi(x)}\,\frac{d\widetilde{\phi}}{\sqrt{2V(\widetilde{\phi})}}\,.$ (5)
@@
Figure 2: The profile of a typical one-dimensional static kink soliton (7) (pictured here for indicative values $m=\sqrt{\lambda}=1$, and $x_{0}=1$). The same profile characterizes the traveling kink wave (10), obtained upon the replacement of $x-x_{0}$ in (7) by the Lorentz boosted coordinate $\gamma\,(x-x_{0}-vt)$, $\gamma=\frac{1}{\sqrt{1-v^{2}}}$, where $v>0\,(<0)$ corresponds to left (right) movers.
@@
[…]
@@
[…]
@@
[…]
@@
The above relation (I) implies that a constant $\phi$ solution is, as expected, characterised by a trivial $Q=0$, but the kink soliton solution is characterised by a non-trivial charge $Q=+1$ ($Q=-1)$ for the anti-kink).
@@
The structure of this article is the following: in the next section II, we review the main theoretical pseudospin model of MT [7], which admits various solitonic solutions, among which helicoidal snoidal waves, which will play an important rôle in our analysis, proving a crucial feature for the rôle of MT as quantum (bio)computers. Such a classical pseudispin model is viewed here as a result of a (partial) collapse of a the newtwork of (quantum) tubulin dimr states, as we explain in some detail. In section III we review the rôle of networks of MTs as classical logic gates, giving emphasis on the rôle of solitons, as well as the Microtubule associated proteins (MAP), which provide a crucial connection across different MTs in the newtwork, thus enforcing their rôle as logic gates. In section IV we describe the basic ingredients for the MTs to operate as scalable biocomputers, namely we discuss the basic information storage unit, the relevant quantum-decoherence mechanism, and the “decision-making” process on the most effective path to be followed for a dissipation-free signal and information transduction across the MT. All the above processes take place within the quantum-decoherence time. We provide a microscopic mechanism, within the QED Cavity model of MT [3, 4, 1], which ensures a relatively long decoherence-time interval, allowing for crucial biocomputing processes to take place. The important rôle of the ordered-water interia of the MT in this respect is highlighted. In section V we discuss an experimental-verification path to be followed in order to falsify or, hopefully, verify (!), the above model, thus supporting further the assumption on the potential rôle of MT as quantum (bio)computers. Specifically, we discuss the Rabi splitting phenomenon, which is associated with the rôle of the entire MT as a cavity, and, if verified, would be a strong indication in favour of the model. We also describe experimental arrangements, and the pertinent measurements, that would probe quantum coherencce and environmental entanglement in individual tubulin dimers, which are the important building blocks of a MT. For the benefit of the reader we summarise our assumptions on the various parameters of the models, and the pertinent physiological conditions, in two tables. Finally, our conclusions and outlook are given in section VI.
@@ II Solitonic effects in Microtubules and observable biological functions
Figure 3: Left : A microtubule (MT) (a), showing individual dimer subunits and their dimensions (b) (1 Angstrom = 0.1 nm). The walls consist of tubulin protein dimers ((c) GTP tubulin, (d) GDP tubulin), which are arranged usually in 12 or 13 helical protofilaments (vertical chain-like structures, parallel to the long axis of MT). The interior (e) is full of ordered water molecules. In the cavity model of MT [3, 1], a thin interior layer near the dimer walls (f) behaves as a high-Q electromagnetic cavity. Right Picture: A network of microtubules typical of the neuronal cytoskeleton. The “rungs” cross-connecting MTs are microtubule associated proteins (MAP [17]) (figures from ref. [1])
Basic features of these models are briefly reviewed here in a modern context of quantum computation in connection with biological systems [27, 28, 29, 30]. Microenvironmental conditions, for in vitro or in vivo MT systems, are found to be crucial for the maintainance of the coherence of quantum effects, when the temperature of the system is within physiological parameters.
@@
In this respect, in ref. [7], we have discussed the emergence of, and classified, various kinds of solitonic excitations, which arise as solutions of the field equations of appropriate classical field theoretic non-linear systems that model the MT dipole interactions in the tubulin dimer. We have used the full cylindrical geometry of MT, but we did not take into account the complicated ordered-water interior, whose role as mentioned above, was assumed simply to provide -through its dipole interactions with the tubulin dipoles - the long decoherence time (14). So although important in this latter respect, the role of the ordered water was not considered further in [7], where we constructed a classical non-linear pseudospin $\sigma$-model, which we assumed to describe the coherent (classically behaving) state of the tubulin dipole quanta after the elapse of the decoherence time (14) (see fig. 4). As discussed in the relevant literature [35], in the ground state of an MT, the orientation of the permanent dipole moments of the tubulin dipoles with respect to the surface of the MT is such that the relative spherical polar angle with the symmetry axis of the MT is $\Theta_{0}\approx 29^{\,\rm o}$. We suggest here that, during the decoherence time (14) (or even in shorter time intervals), in analogy with the situation in Algae, the MT “quantum computes” the most efficient pathway for energy transfer, which is realized by the formation of the appropriate solitons, in the way we describe below.
@@
$J_{ij}=\frac{1}{4\pi\epsilon\,\epsilon_{0}\,r_{ij}^{3}}\,,$ (16)
@@
where $\epsilon$ is the permittivity of the MT microenvironment in units of that of the vacuum, $\epsilon_{0}$, and $r_{ij}$ is the distance between sites $i$ and $j$ of the lattice model. As is common, in [7] we assumed that $J_{ij}$, are nonzero only for the nearest-neighbor dipole moments (in practice, next to nearest neighbor dipole-dipole interactions are considered suppressed). The middle term, with $P$ and $Q$ appropriate interaction couplings, has the form of a double-well quartic on-site potential, and takes into account [39] the assumed ferroelectric properties at physiological temperature ranges for the MT [40, 35], and their effects on the effective spin, $\mathbf{S}_{i}$, while the last term , describes the effects of the transversal (radial) electrostatic field with amplitude $B$ acting on the dipoles, which is produced by the solvent environment of the MT. Thus, all the further effects of the ordered water molecules (apart from their important contribution to lead to the long decoherence time (14)), are captured by this term.
@@
The system of MT dimers may be represented as a triangular lattice, as shown in Fig. 5, so that each spin has six nearest neighbors. The constants of interaction between the central dipole (labelled “00”) in Fig. 5 and its nearest neighbors are denoted as $J_{0\alpha}$, and the distance between the central spin and its nearest neighbors as $d_{\alpha}$ ($\alpha=1,2,\dots,6$). We set $d_{01}=d_{04}=a$, $d_{02}=d_{05}=b$, $d_{03}=d_{06}=c$. The corresponding angles (between the central dimer and others) are denoted as, $\theta_{1}$, $\theta_{2}$ and $\theta_{3}$, so that: ${\mathbf{e}}_{01}\cdot{\mathbf{e}}_{01}=\cos\theta_{1}$, ${\mathbf{e}}_{01}\cdot{\mathbf{e}}_{02}=\cos\theta_{2}$, ${\mathbf{e}}_{01}\cdot{\mathbf{e}}_{06}=\cos\theta_{3}$. Typical values of parameters known from the literature are: $a=8\,\rm nm$, $b=5.87\,\rm nm$, $c=7.02\,\rm nm$, $\theta_{1}=0$, $\theta_{2}=58.2^{\,\rm o}$, $\theta_{3}=45.58^{\,\rm o}$, $S=1714$ Debye  [41, 39] (See Fig. 5b.) The radius of the MT can be estimated as, $R\approx 11.2\,\rm nm$ [35, 42]. The unit cell shown in Fig. 5 consists of the central spin surrounded by six neighbors. Its area is: $\Sigma_{0}=3ad=120\,\rm nm^{2}$.
@@
Figure 5: Tubulin neighborhood in the hexagonal unit cell of the microtubule. The distance between dimers is $d$. The heterodimer helix direction is defined by the height, $h$. The typical values of parameters are: $a=8\,\rm nm$, $b=5.87\,\rm nm$, $c=7.02\,\rm nm$, $d=5\,\rm nm$, $h=4.9\,\rm nm$, $\theta_{1}=0$, $\theta_{2}=58.2^{\,\rm o}$, $\theta_{3}=45.58^{\,\rm o}$ [41, 35, 39, 42]. This structure will play the rôle of the quDit basic information storage unit in our modelling of the MT as a biocomoputer.
@@
As discussed in [7], the continuum approximation proved sufficient for classifying and studying the solitonic solutions arising from the non-linear lagrangian corresponding to the pseudospin non-linear $\sigma$-model with interaction potential given by (15). Using the local spherical coordinates $(\Theta_{i},\Phi_{i})$ to define the orientation of the dipole, $\mathbf{n}=(\sin\Theta\cos\Phi,\sin\Theta\sin\Phi,\cos\Theta)$, then, the continuum Lagrangian of the system reads:
@@
with $h=(6S^{2}/J)\,\sum_{a=1}^{3}J_{0a}\,{\rm cos}^{2}\theta_{a}$, $g_{0}=P\,S^{2}/J,\,g_{1}=Q\,S^{4}/J,\,g_{2}=B\,S/J$, and $J=2\,S^{2}\,\sum_{a}^{3}J_{0a}$ On introducing the dimensionless coordinates, $\zeta=z/\sqrt{\Sigma_{0}}$ and $\tilde{R}=R/\sqrt{\Sigma_{0}}$, the continuum Lagrangian of the system becomes that of an anisotropic $\sigma$-model:
@@
In [7] the zero-temperature phase diagram of this model has been studied in detail, yielding for the ground state (which is characterized by a permanent dipole moment) a paraelectric and a ferroelectric phase, separated by the line $\kappa=4\sigma$, where $\sigma\equiv=(h-g_{0})/(2g_{1})$ and $\kappa\equiv g_{2}/g_{1}$. The ferroelectric phase of the MT ground state occurs for $\sigma>0$ and $\kappa=g_{2}/g_{1}<4\,\sigma$, while the paraelectric phase occurs in the regime of parameters $\sigma<0$, and $\kappa>4\sigma$. This zero-temperature paraelectric phase corresponds to the radial orientation of the permanent dipole moments of the tubulin dimers with respect to the surface of the MT. For finite temperatures of interest to realistic MT systems, we refer the reader to [43] where the critical order-disorder transition temperature depends on the values of the dipole moment and the electric permittivity of the system. It will not be of further impact to our considerations in this paper. We only remark for completion that a discussion on the importance of ferroelectricity in biological systems has been given in [44], and it has been at the heart of the concrete MT modelling since the early days [32], [3].
@@
where we set $\sigma_{0}^{2}=1-1/h$. For completeness, we mention that the analysis of [7], taking into account the parameters entering the model of MT under consideration here, shows that dependence of $v$ on the parameter that the velocity of the wave is bounded from above $v\leq v_{0}$, where $v_{0}\approx 155\rm m/s$.
@@
[…]

### v2

@@ Abstract
We examine the quantum coherence properties of tubulin heterodimers arranged into the protofilaments of cytoskeletal microtubules. In the physical model proposed by the authors, the microtubule interiors are treated as high-Q quantum electrodynamics (QED) cavities that can support decoherence-resistant entangled states under physiological conditions, with decoherence times of the order of $\mathcal{O}(10^{-6})$ sec. We identify strong electric dipole interactions between tubulin dimers and ordered water dipole quanta within the microtuble interior as the mechanism responsible for the extended coherence times. Classical nonlinear (pseudospin) $\sigma$-models describing solitonic excitations are reinterpreted as emergent quantum-coherent—or possibly pointer—states, arising from incomplete collapse of dipole-aligned quantum states. These solitons mediate dissipation-free energy transfer along microtubule filaments. We discuss logic-gate-like behavior facilitated by microtubule-associated proteins, and outline how such structures may enable scalable, ambient-temperature quantum computation, with the fundamental unit of information storage realized as a quDit encoded in the tubulin dipole state. We further describe a process akin to ”decision making” that emerges following an external stimulus, whereby optimal, energy-loss-free signal and information transport pathways are selected across the microtubular network. Finally, we propose experimental approaches—including Rabi-splitting spectroscopy and entangled surface plasmon probes—to validate the use of biomatter as a substrate for scalable quantum computation.
@@ I Introduction: Solitons, dissipationless energy transfer and quantum entanglement in biosystems
The role of biomatter in quantum computation remains an open question. This issue is explored in [1] through a model of cytoskeletal microtubules (MTs) [2], treated as high-quality quantum electrodynamics (QED) cavities [3, 4], with a particular focus on the efficiency of energy—and therefore information—transduction and transport in biological systems. In our previous work, we argued that the QED cavity model of MTs may be instrumental in understanding memory encoding and retrieval functions [5], as it enables the conversion of external stimulus into corresponding electromagnetic signals. QED cavities are well known for facilitating efficient manipulation of quantum entanglement between atoms and photons [6]. In an analogous manner, we proposed in [3] that the electric dipole quanta associated with tubulin dimer walls in the MT cavity interior can become entangled with the dipole quanta of the ordered water in the MT core, giving rise to solitonic dipole structures that mediate dissipation-free energy and signal transport.
Microtubule (MT) dipole systems can be modeled using condensed-matter approaches, such as pseudospin nonlinear $\sigma$-models [7], allowing loss-free signal transduction to be quantified via explicit solitonic solutions to the corresponding equations of motion. As shown in [3, 4, 1] and expanded upon here, these classical solitons may be viewed as coherent states of the underlying dipole quanta. [^1] MTs are ubiquitous structures in all eukaryotic cells [2], but there are many other biomolecules involved in memory encoding and transmission for instance: recently it has been proved possible to transplant memories affecting behavior carried by RNA molecules (which themselves have been known to carry genetic information for decades  [8], but not actual cognitive memories). Direct learned memory transfer by RNA injection has been shown between snails  [9], that is, passing down RNA through flesh creates memories in the recipient instead of the conventional path of perceptions translated into neurotransmitters and action potentials through an intact nervous system. Our model for information processing may generally apply to such biological structures as well but treatment of them is outside the scope of this paper.
@@
In the 1990’s, a suggestion on the role of quantum effects on brain functioning, and in particular on conscious perception, has been put forward by S. Hameroff and R. Penrose (HP) [15], who concentrated on MTs [2] of the brain cells. Specifically, by considering the tubulin heterodimers conformation as quantum states of a two-state system, they assumed their coherent superposition, which may result in helicoidal solitonic states propagating along the MT, [^3] Note that the helix is a geodesic path (optimal “minimum distance” path) in the Euclidean cylindrical geometry $S^{1}\otimes\mathbb{R}$, embeddable in $\mathbb{R}^{3}$. and thus being responsible for conscious perception. HP assumed sufficiently long decoherence time for this purpose of order ${\mathcal{O}}(1~s)$ in in vivo situations, so that the in vivo system of MT in the brain undergoes, as a result of sufficient growth that allowed it to reach a critical mass/energy, self-collapse related to a quantum gravity environment (orchestrated reduction method), as opposed to the standard environmental non-gravitational decoherence that physical quantum systems are subjected to [13].
@@
$x-x_{0}$ $=\pm\int_{\phi(x_{0})}^{\phi(x)}\,\frac{d\widetilde{\phi}}{\sqrt{2V(\widetilde{\phi})}}\,$. (5)
@@
Figure 2: The profile of a typical one-dimensional static kink soliton (7) (pictured here for indicative values $m=\sqrt{\lambda}=1$, and $x_{0}=1$). The same profile characterizes the traveling kink wave (10), obtained upon the replacement of $x-x_{0}$ in (7) by the Lorentz boosted coordinate $\gamma\,(x-x_{0}-vt)$, $\gamma=\frac{1}{\sqrt{1-v^{2}}}$, where v>0(<0)v>0\,(<0) corresponds to left (right) movers.
@@
$M=\int_{-\infty}^{\infty}\mathcal{E}(x)\,dx=\frac{2\sqrt{2}}{3}\,\frac{m^{3}}{\lambda}\,$, (9)
@@
The way the solitons behave at spatial infinity (x = $\pm\infty$ ) imply that it takes an infinite amount of energy to change the (anti)kink configuration. This is an indication of its stability, but the latter is also a guarantee -rigorously for topological reasons. Indeed, for the system (1), although there are no Noether currents associated with a symmetry of the real scalar field, nevertheless, there is a (trivially) conserved current
@@
$Q$ $=\int_{-\infty}^{\infty}dx\mathcal{J}^{0}=\frac{1}{2}\,\frac{\sqrt{\lambda}}{m}\int_{-\infty}^{\infty}dx\,\phi^{\prime}(x)$
@@
The above relation (I) implies that a constant $\phi$ solution is, as expected, characterised by a trivial $Q=0$, but the kink soliton solution is characterised by a non-trivial charge $Q=+1$ (Q=1)Q=-1) for the anti-kink).
@@
The structure of this article is the following: in the next section II, we review the main theoretical pseudospin model of MT [7], which admits various solitonic solutions, among which helicoidal snoidal waves, which will play an important rôle in our analysis, proving a crucial feature for the rôle of MT as quantum (bio)computers. Such a classical pseudispin model is viewed here as a result of a (partial) collapse of a the newtwork of (quantum) tubulin dimr states, as we explain in some detail. In section III we review the rôle of networks of MTs as classical logic gates, giving emphasis on the rôle of solitons, as well as the Microtubule associated proteins (MAP), which provide a crucial connection across different MTs in the newtwork, thus enforcing their rôle as logic gates. In section IV we describe the basic ingredients for the MTs to operate as scalable biocomputers, namely we discuss the basic information storage unit, the relevant quantum-decoherence mechanism, and the “decision-making” process on the most effective path to be followed for a dissipation-free signal and information transduction across the MT. All the above processes take place within the quantum-decoherence time. We provide a microscopic mechanism, within the QED Cavity model of MT [3, 4, 1], which ensures a relatively long decoherence-time interval, allowing for crucial biocomputing processes to take place. The important rôle of the ordered-water interia of the MT in this respect is highlighted. In section V we discuss an experimental-verification path to be followed in order to falsify or, hopefully, verify (!), the above model, thus supporting further the assumption on the potential rôle of MT as quantum (bio)computers. Specifically, we discuss the Rabi splitting phenomenon, which is associated with the rôle of the entire MT as a cavity, and, if verified, would be a strong indication in favour of the model. We also describe experimental arrangements, and the pertinent measurements, that would probe quantum coherence and environmental entanglement in individual tubulin dimers, which are the important building blocks of a MT. For the benefit of the reader we summarise our assumptions on the various parameters of the models, and the pertinent physiological conditions, in two tables. Finally, our conclusions and outlook are given in section VI.
@@ II Solitonic effects in Microtubules and observable biological functions
Figure 3: Left : A microtubule (MT) (a), showing individual dimer subunits and their dimensions (b) (1 Angstrom = 0.1 nm). The walls consist of tubulin protein dimers ((c) GTP tubulin, (d) GDP tubulin), which are arranged usually in 12 or 13 helical protofilaments (vertical chain-like structures, parallel to the long axis of MT). The interior (e) is full of ordered water molecules. In the cavity model of MT![3, 1], a thin interior layer near the dimer walls (f) behaves as a high-Q electromagnetic cavity. Right Picture: A network of microtubules typical of the neuronal cytoskeleton. The “rungs” cross-connecting MTs are microtubule associated proteins (MAP [17]) (figures from ref. [1])
Basic features of these models are briefly reviewed here in a modern context of quantum computation in connection with biological systems [27, 28, 29, 30]. Microenvironmental conditions, for in vitro or in vivo MT systems, are found to be crucial for the maintainance of the coherence of quantum effects, when the temperature of the system is within physiological parameters. For the convenience of the reader we give in Table 1 a glossary of symbols to be used in what follows, together with their meanings (equation numbers refer to the manuscript).
Table 1: Glossary of symbols (equation numbers refer to the manuscript).
Symbol Meaning First defined / used Units
$\varepsilon$ Relative dielectric constant (MT interior/medium) Eq. (34), Table 4
$E_{\mathrm{ow}}$ r.m.s. ordered-water field amplitude Eq.  (34) V m-1
$\omega_{0}$ Tubulin dimer transition frequency Sec. V.1 Eq. (32) s-1
$\omega_{c}$ Dominant cavity mode frequency Eq. (35) s-1
$\Delta$ Detuning, $\Delta=\omega_{0}-\omega_{c}$ Eq. (32) s-1
$\lambda_{0}$ Single-dimer vacuum Rabi coupling, $dE_{\rm ow}/\hbar$ Eq. (33) s-1
NN Number of dimers in the cavity mode Eq. (38)
$\Omega_{\pm}$ Absorption peaks (Rabi branches) Eq. (32) s-1
$t_{\mathrm{decoh}}$ Decoherence time (cavity-limited) Eq. (14) s
$J_{ij}$ Dipole–dipole coupling Eq. (16) energy
$h,g_{0},g_{1},g_{2}$ Pseudospin $\sigma$-model parameters Eqs.  (17)–(19)
$\kappa$ $g_{2}/g_{1}$ (phase-diagram parameter) Sec. II (below Eq.  (19))
$\sigma$ $(h-g_{0})/(2g_{1})$ (phase-diagram parameter) Sec. II (below Eq. (19))
$\Sigma_{0}$ Area of hexagonal unit cell Fig. 5 nm2
$\rho$ Areal mass density (continuum limit) Eq.  (17) context kg m-2
$u_{0},\ \zeta_{0}$ Velocity parameters ($\zeta_{0}^{2}=1-1/h$) Eq. (22)
$k$ Elliptic modulus of snoidal solutions Eqs. (23)–(25), Fig. 7
$\xi$ Soliton solution argument Eq. (20)
@@
In this respect, in ref. [7], we have discussed the emergence of, and classified, various kinds of solitonic excitations, which arise as solutions of the field equations of appropriate classical field theoretic non-linear systems that model the MT dipole interactions in the tubulin dimer. We have used the full cylindrical geometry of MT, but we did not take into account the complicated ordered-water interior, whose role as mentioned above, was assumed simply to provide -through its dipole interactions with the tubulin dipoles - the long decoherence time (14). So although important in this latter respect, the role of the ordered water was not considered further in [7], where we constructed a classical non-linear pseudospin $\sigma$-model, which we assumed to describe the coherent (classically behaving) state of the tubulin dipole quanta after the elapse of the decoherence time (14) (see fig. 4). As discussed in the relevant literature [35], in the ground state of an MT, the orientation of the permanent dipole moments of the tubulin dipoles with respect to the surface of the MT is such that the relative spherical polar angle with the symmetry axis of the MT is $\Theta_{0}\approx 29^{\,\rm o}$. We suggest here that, during the decoherence time (14) (or even in shorter time intervals), in analogy with the situation in Algae, the MT ‘quantum computes’ the most efficient pathway for energy transfer, which is realized by the formation of the appropriate solitons, in the way we describe below.
@@
$J_{ij}=\frac{1}{4\pi\varepsilon\,\epsilon_{0}\,r_{ij}^{3}}\,$, (16)
@@
where $\varepsilon$ is the permittivity of the MT microenvironment in units of that of the vacuum, $\epsilon_{0}$, and $r_{ij}$ is the distance between sites $i$ and $j$ of the lattice model. As is common, in [7] we assumed that $J_{ij}$, are nonzero only for the nearest-neighbor dipole moments (in practice, next to nearest neighbor dipole-dipole interactions are considered suppressed). The middle term, with $P$ and $Q$ appropriate interaction couplings, has the form of a double-well quartic on-site potential, and takes into account [39] the assumed ferroelectric properties at physiological temperature ranges for the MT [40, 35], and their effects on the effective spin, $\mathbf{S}_{i}$, while the last term , describes the effects of the transversal (radial) electrostatic field with amplitude $B$ acting on the dipoles, which is produced by the solvent environment of the MT. Thus, all the further effects of the ordered water molecules (apart from their important contribution to lead to the long decoherence time (14)), are captured by this term.
@@
The system of MT dimers may be represented as a triangular lattice, as shown in Fig. 5, so that each spin has six nearest neighbors. The constants of interaction between the central dipole (labelled “0”) in Fig. 5 and its nearest neighbors are denoted as $J_{0\alpha}$, and the distance between the central spin and its nearest neighbors as $d_{\alpha}$ ($\alpha=1,2,\dots,6$). We set $d_{01}=d_{04}=a$, $d_{02}=d_{05}=b$, $d_{03}=d_{06}=c$. The corresponding angles (between the central dimer and others) are denoted as, $\theta_{1}$, $\theta_{2}$ and $\theta_{3}$, so that: ${\mathbf{e}}_{01}\cdot{\mathbf{e}}_{01}=\cos\theta_{1}$, ${\mathbf{e}}_{01}\cdot{\mathbf{e}}_{02}=\cos\theta_{2}$, ${\mathbf{e}}_{01}\cdot{\mathbf{e}}_{06}=\cos\theta_{3}$. Typical values of parameters known from the literature are: $a=8\,\rm nm$, $b=5.87\,\rm nm$, $c=7.02\,\rm nm$, $\theta_{1}=0$, $\theta_{2}=58.2^{\,\rm o}$, $\theta_{3}=45.58^{\,\rm o}$, $S=1714$ Debye  [41, 39] (See Fig. 5b.) The radius of the MT can be estimated as, $R\approx 11.2\,\rm nm$ [35, 42]. The unit cell shown in Fig. 5 consists of the central spin surrounded by six neighbors. Its area is: $\Sigma_{0}=3ad=120\,\rm nm^{2}$.
@@
Figure 5: Tubulin neighborhood in the hexagonal unit cell of the microtubule. The distance between dimers is $d$. The heterodimer helix direction is defined by the height, $h$. The typical values of parameters are: $a=8\,\rm nm$, $b=5.87\,\rm nm$, $c=7.02\,\rm nm$, $d=5\,\rm nm$, $h=4.9\,\rm nm$, $\theta_{1}=0$, $\theta_{2}=58.2^{\,\rm o}$, $\theta_{3}=45.58^{\,\rm o}$ [41, 35, 39, 42]. This structure will play the rôle of the quDit basic information storage unit in our modelling of the MT as a biocomputer.
@@
As discussed in [7], the continuum approximation proved sufficient for classifying and studying the solitonic solutions arising from the non-linear lagrangian corresponding to the pseudospin non-linear $\sigma$-model with interaction potential given by (15). Using the local spherical coordinates $(\Theta_{i},\Phi_{i})$ to define the orientation of the dipole, $\mathbf{n}=(\sin\Theta\cos\Phi,\sin\Theta\sin\Phi,\cos\Theta)$, the continuum Lagrangian of the system reads:
@@
with $h=(6S^{2}/J)\,\sum_{a=1}^{3}J_{0a}\,{\rm cos}^{2}\theta_{a}$, $g_{0}=P\,S^{2}/J,\,g_{1}=Q\,S^{4}/J,\,g_{2}=B\,S/J$, and $J=2\,S^{2}\,\sum_{a}^{3}J_{0a}$ . On introducing the dimensionless coordinates, $\zeta=z/\sqrt{\Sigma_{0}}$ and $\tilde{R}=R/\sqrt{\Sigma_{0}}$, the continuum Lagrangian of the system becomes that of an anisotropic $\sigma$-model:
@@
In [7] the zero-temperature phase diagram of this model has been studied in detail, yielding for the ground state (which is characterized by a permanent dipole moment) a paraelectric and a ferroelectric phase, separated by the line $\kappa=4\sigma$, where $\sigma\equiv(h-g_{0})/(2g_{1})$ and $\kappa\equiv g_{2}/g_{1}$. The ferroelectric phase of the MT ground state occurs for $\sigma>0$ and $\kappa=g_{2}/g_{1}<4\,\sigma$, while the paraelectric phase occurs in the regime of parameters $\sigma<0$, and $\kappa>4\sigma$. This zero-temperature paraelectric phase corresponds to the radial orientation of the permanent dipole moments of the tubulin dimers with respect to the surface of the MT. For finite temperatures of interest to realistic MT systems, we refer the reader to [43] where the critical order-disorder transition temperature depends on the values of the dipole moment and the electric permittivity of the system. It will not be of further impact to our considerations in this paper. We only remark for completion that a discussion on the importance of ferroelectricity in biological systems has been given in [44], and it has been at the heart of the concrete MT modelling since the early days [32], [3].
@@
where we set $\sigma_{0}^{2}=1-1/h$. For completeness, we mention that the analysis of [7], taking into account the parameters entering the model of MT under consideration here, shows that the velocity of the wave is bounded from above $v\leq v_{0}$, where $v_{0}\approx 155\rm m/s$ (see figure 6).
Figure 6: The (bounded) velocity (22) (in m/s) of the non-linear wave as a function of the variable $u_{0}$, for two different values of the tubulin dimer mass (viewed as a dipole of length $l$) $M=10^{-22}\,\rm g$ (red) and $M=10^{-23}\,\rm g$ (blue), with $l\simeq 2$ nm. Picture taken from [7].
@@
with the boundary conditions: $u(\pm\infty)=\pm 1$. [^5] Kinks, as is well known [10, 46], and discussed in the introduction of this article, admit a topological classification in terms of the appropriate homotopy group. In our case, the topological charge, $\pi_{0}$, of the kink (7) is determined by the magnitude, $n_{z}$ of the polarization vector at the ends of the MT: $\pi_{0}=\frac{1}{2}(n_{z}(+\infty)-n_{z}(-\infty)).$ (26) To change the topological charge one needs to overcome the potential barrier, proportional to the size of the MT (formally, infinite potential barrier).